To generate tension in a muscle, a myosin molecule pulls on an actin filament with a force of 1.50 µN. if the actin filament moves at 3.56 nm/s, what is the power output P of this molecular motor? Use prefixes to express your answer.


1
Expert's answer
2021-06-29T05:14:02-0400

Power in terms of Force and velocity is given by

"P=F\times v\\
F=1.5\mu N=1.5\times 10^{-6}N\\
v=3.56nm/s\\
v=3.56\times 10^{-9}m/s\\
P=1.5\times 10^{-6}\times3.56\times 10^{-9}\\
P=3.56fW"